OpenStudy (lexy148):

An investment grows at a rate of 6% per year. How much, to the nearest $100, should Noel invest if he wants to have $100,000 at the end of 20 years?

OpenStudy (photon336):

A = P*r*t where A is the final value of the investment P = the starting value, r = the rate in decimal form and t = time in years.

OpenStudy (photon336):

We need to figure out how much he has to invest initially so we solve the equation for P the starting value. then you just plug in your values. \[\frac{ A }{ r*T } = P \]

OpenStudy (lexy148):

yes i understand how to do these problems, the problem is idk how to solve it. i was thinking of it being A=100,000(1+.06)^20 but im not sure if thats the correct way to do it

OpenStudy (photon336):

you need to find your original value invested

OpenStudy (lexy148):

oh okay, so A would equal to 100,000?

OpenStudy (photon336):

Is the interest compounded?

OpenStudy (lexy148):

no it isnt compounded

satellite73 (satellite73):

solve \[P(1.06)^{20}=1000\] for \(P\)

satellite73 (satellite73):

with a calculator, compute \[P=\frac{100,000}{(1.06)^{20}}\]

OpenStudy (photon336):

Your final value should be \[\frac{ 100,000 }{ (1.06)^{20} }\]

OpenStudy (photon336):

i'm wondering though why we can't also just use simple interest since it's not compounded @satellite73

OpenStudy (lexy148):

okay so my answer should be 31,180

satellite73 (satellite73):

it says 6% per year, all interest is compounded